can you peove this?:cosbcscbtanb=1
Answer by ad_alta(240)

Let 'o' be opposite, 'a' adjacent, and 'h' hypotenuse. We can rewrite the expression as (a/h)(h/o)(o/a)=(a/o)(o/a)=1. This actually isn't always true, so it isn't really an "identity." There is a strong requirement that h, o, and a be nonzero, which is certainly not true for all b.